### v1

# Making Parametric Anomaly Detection on Tabular Data Non-Parametric Again

Abstract.
Deep learning for tabular data has garnered increasing attention in recent years, yet employing deep models for structured data remains challenging. While these models excel with unstructured data, their efficacy with structured data has been limited. Recent research has introduced retrieval-augmented models to address this gap, demonstrating promising results in supervised tasks such as classification and regression. In this work, we investigate using retrieval-augmented models for anomaly detection on tabular data. We propose a reconstruction-based approach in which a transformer model learns to reconstruct masked features of normal samples. We test the effectiveness of KNN-based and attention-based modules to select relevant samples to help in the reconstruction process of the target sample. Our experiments on a benchmark of 31 tabular datasets reveal that augmenting this reconstruction-based anomaly detection (AD) method with non-parametric relationships via retrieval modules may significantly boost performance. Our code to reproduce the experiments is made available on GitHub.
Anomaly Detection, Tabular Data, Deep Learning
1. Introduction
Supervised anomaly detection (AD) consists in learning to characterize a normal distribution using a dataset only composed of samples belonging to the normal [^1] The term normal relates to the concept of normality, in opposition to anomaly. class, in order to identify in a separate dataset the samples that do not belong to this normal distribution, namely anomalies. This type of approach is often used when the imbalance between classes is too severe, causing standard supervised approaches to fail (Yanmin et al., 2011). Examples of such applications are cyber intrusion detection (Ben Said et al., 2023), fraud detection on credit card payment (Hilal et al., 2022; Thimonier et al., 2023b), or tumor detection on images (Wei et al., 2018). On the contrary, Unsupervised anomaly detection refers to identifying anomalies in a dataset without using labeled training data. These algorithms aim to discover patterns or structures in the data and flag instances that deviate significantly from these patterns. The application of such an approach usually includes detecting mislabeled samples or removing outliers from a dataset that may hinder the training of a model.
While deep learning methods have become ubiquitous and are widely used in the industry for various tasks on unstructured data, relying on deep models for tabular data remains challenging. Indeed, Grinsztajn et al. (2022) discuss how the inherent characteristics of tabular data make this type of data challenging to handle by standard deep models. Hence, recent research on deep learning for structured data (Shavitt and Segal, 2018; Arik and Pfister, 2021; Somepalli et al., 2021; Kossen et al., 2021; Gorishniy et al., 2023) has been oriented towards proposing novel training frameworks, regularization techniques or architectures, tailored for tabular data. Similarly, the best-performing anomaly detection algorithms proposed in the literature involve accounting for the particular structure of tabular data while general approaches tend to struggle. For instance, in (Bergman and Hoshen, 2020; Qiu et al., 2021; Shenkar and Wolf, 2022; Thimonier et al., 2023a) authors put forward self-supervised anomaly detection algorithms targeted for tabular data.
In particular, recent research has emphasized the pivotal role of non-parametric relations between samples in fostering deep learning model’s performance (Gorishniy et al., 2023; Kossen et al., 2021; Somepalli et al., 2021; Thimonier et al., 2023a). Kossen et al. (2021) and Somepalli et al. (2021) propose novel deep architecture with internal retrieval models that allow leveraging inter-sample relations. Similarly, Gorishniy et al. (2023) investigate the pertinence of including non-parametric mechanisms by augmenting an MLP with external retrieval modules and show a significant increase in performance for classification and regression tasks on tabular data. Parallel to this line of work, the reconstruction-based AD method of Thimonier et al. (2023a) relies on non-parametric transformers (Kossen et al., 2021) to leverage sample-sample dependencies and shows significant improvement in anomaly detection accuracy in comparison with existing methods.
As empirically demonstrated in (Thimonier et al., 2023a), both feature-feature and sample-sample dependencies should be conjointly learned by a model to detect anomalies on tabular data efficiently. Indeed, Han et al. (2022) discusses how anomalies on tabular data can be categorized into $4$ types of anomalies. We believe leveraging both interactions is necessary to efficiently identify all $4$ types. First, dependency anomalies, referring to samples that do not follow the dependency structure that normal data follow, require feature-feature dependencies to be efficiently identified. Second, global anomalies refer to unusual data points that deviate significantly from the norm. Relying on both dependencies should improve a model’s capacity to detect these anomalies. Third, local anomalies that refer to the anomalies that are deviant from their local neighborhood can only be identified by relying on sample-sample dependencies. Finally, clustered anomalies, also known as group anomalies, are composed of anomalies that exhibit similar characteristics. This type of anomaly requires feature-feature dependencies to be identified.
Hence, we further explore the possibility of leveraging both feature-feature and sample-sample relations for supervised anomaly detection by investigating how existing AD methods can be augmented by an external retrieval module. External retrieval modules can be considered as instrumental as they can augment any existing model. In contrast, most models that rely on internal retrieval mechanisms (e.g. non-parametric transformers (Kossen et al., 2021)) are bound to some inductive biases and cannot be used to learn all possible tasks that may be relevant for anomaly detection. We test the relevance of external modules by employing transformers (Vaswani et al., 2017) in a mask-reconstruction framework to construct an anomaly score since it was proven to offer strong anomaly detection performance (Thimonier et al., 2023a). We implement several external retrieval methods to augment the vanilla transformer and evaluate the performance of each approach on an extensive benchmark of tabular datasets. We empirically show that the tested approaches incorporating retrieval modules to account for the sample-sample relations outperform the vanilla transformer that only attends to feature-feature dependencies.
The present work offers the following contributions:
  • We propose an extensive evaluation of retrieval-based methods for AD on tabular data.
  • We empirically show that augmenting existing AD methods with a retrieval module to leverage sample-sample dependencies can help improve detection performance.
2. Related works
Ruff et al. (2021) discuss how anomaly detection bears several denominations that more or less designate the same class of algorithms: anomaly detection, novelty detection, and outlier detection. Overall, the literature is comprised of $4$ main classes of anomaly detection algorithms: density estimation, one-class classification, reconstruction-based and self-supervised algorithms.
Density estimation
It is often seen as the most direct approach to detecting anomalies in a dataset. The density estimation approach consists in estimating the normal distribution and flagging low probability samples under this distribution as an anomaly. Existing methods include using Copula as the COPOD method proposed in (Li et al., 2020), Local Outlier Factor (LOF) (Breunig et al., 2000), Energy-based models (Zhai et al., 2016) flow-based models (Liu et al., 2022).
Reconstruction-based methods
Other anomaly detection methods focus on learning to reconstruct samples belonging to the normal distribution. In inference, the capacity of the model to reconstruct an unseen sample is used as a measure of anomalousness. The more capable a model is to reconstruct a sample, the more likely the sample is to belong to the normal distribution seen in training. Such approach include methods involving autoencoders (Chen and Konukoglu, 2018; Kim et al., 2020), diffusion models (Wolleb et al., 2022; Zhang et al., 2023), GANs (Schlegl et al., 2017) or attention-based models (Thimonier et al., 2023a).
One-Class Classification
One-class classification describes the task of identifying anomalies without directly estimating the normal density. This class of algorithm involves discriminative models that directly estimate a decision boundary. In (Schölkopf et al., 1999; Tax and Duin, 2004; Ruff et al., 2018), the authors propose algorithms that estimate the support of the normal distribution, either in the original data space or in a latent space. During inference, one flags the samples outside the estimated support as anomalies. Other one-class classification methods include tree-based approaches such as isolation forest (IForest) (Liu et al., 2008), extended isolation forest (Hariri et al., 2021), RRCF (Guha et al., 2016) and PIDForest (Gopalan et al., 2019). Other methods also include approaches to augment existing one-class classification methods with a classifier through generating synthetic anomalies during training, such as DROCC (Goyal et al., 2020).
Self-Supervised Approaches
Given the recent successes of self-supervision for a large variety of tasks, researchers have also investigated using self-supervised methods for anomaly detection. Bergman and Hoshen (2020) and Qiu et al. (2021) propose transformation based anomaly detection methods for tabular data. The former relies on a classifier’s capacity to identify which transformation was applied to a sample to measure anomalousness, while the latter relies on a contrastive approach. Similarly, Shenkar and Wolf (2022) also propose a contrastive approach to flag anomalies by learning feature-feature relation for normal samples. Parallel to this line of work, (Sohn et al., 2021; Reiss and Hoshen, 2023) have focused on proposing self-supervised approaches for representation learning tailored for anomaly detection.
Retrieval modules
Retrieval modules have gained attention in recent years in many fields of machine learning. For instance, Blattmann et al. (2022) introduces a retrieval module to foster the scalability and efficiency of diffusion models. Parallel to that, Lin et al. (2022) introduced retrieval for cross-task generalization of large language models, and Chen et al. (2022) introduced it to enhance prompt learning. Finally, retrieval methods have been increasingly used to increase the performance of deep models for tabular data. For instance, Kossen et al. (2021) and Somepalli et al. (2021) introduced internal retrieval modules in deep architecture for supervised tasks on tabular data, while Thimonier et al. (2023a) relied on internal retrieval modules for anomaly detection. Finally, Gorishniy et al. (2023) investigated using external retrieval modules to augment an MLP for supervised tasks on tabular data.
[…]
Figure 1. Forward pass for sample $\mathbf{z}$, see appendix A.2.1 for more detail on training procedure. In the case of no retrieval module, the prediction for a sample $\mathbf{z}$ consists of the upper part of the figure with $\lambda=0$.
3. Method
3.1. Learning Objective
Let $\mathcal{D}_{train}=\{\mathbf{x}_{i}\in\mathbb{R}^{d}\}_{i=1}^{n}$ represent the training set composed of $n$ normal samples with $d$ features. The standard approach to anomaly detection involves learning some function $\phi_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{p}$ by minimizing a loss function $\mathcal{L}$. The chosen loss function will vary according to the class of the considered anomaly detection algorithm. Nevertheless, the overall aim of $\mathcal{L}$ is to characterize the distribution of the samples in the training set as precisely as possible. Depending on the chosen AD algorithm, the obtained representation $\phi_{\theta}(\mathbf{x})$ of sample $\mathbf{x}$ can be used directly or indirectly to compute an anomaly score.
Formally, the training objective can be summarized as follows
(1) $\min_{\theta\in\Theta}\sum_{\mathbf{x}\in\mathcal{D}_{train}}\mathcal{L}(\mathbf{x},\phi_{\theta}(\mathbf{x})),$
where $\mathcal{L}(\mathbf{x},\phi_{\theta}(\mathbf{x}))$ will vary according to the chosen task. For a reconstruction-based method $\mathcal{L}$ can be the squared $\ell_{2}$-norm of the difference between the original sample $\mathbf{x}$ and its reconstructed counterpart, $\mathcal{L}(\mathbf{x},\phi_{\theta}(\mathbf{x}))=\|\mathbf{x}-\phi_{\theta}(\mathbf{x})\|^{2}$.
Introducing an external retrieval module permits keeping the original objective unchanged while augmenting $\phi_{\theta}$ with non-parametric mechanisms. The model involves non-parametric relations because it leverages the entire training dataset when making its prediction. Hence, to optimize its objective the model can conjointly attend to feature-feature and sample-sample interactions.
Formally, instead of minimizing the loss as described in equation 1, the parameters $\theta$ of function $\phi_{\theta}$ are optimized to minimize the loss function as follows
(2) $\min_{\theta\in\Theta}\sum_{\mathbf{x}\in\mathcal{D}_{train}}\mathcal{L}\left(\mathbf{x},\phi_{\theta}\left(\mathbf{x};\mathcal{D}_{train}\right)\right).$
Nevertheless, not all approaches to AD may benefit from such non-parametric mechanisms. Some pretext tasks involving sample-sample dependencies may lead to degenerate solutions, e.g. approaches based on contrastive learning such as the approaches of Qiu et al. (2021) or Shenkar and Wolf (2022). However, reconstruction-based AD methods appear as a natural class of algorithms that may benefit from these non-parametric mechanisms.
Mask Reconstruction
We empirically investigate the pertinence of external retrieval modules, as detailed in section 3.2, on a similar method to Thimonier et al. (2023a) involving stochastic masking and mask reconstruction for tabular data. Stochastic masking consists in masking each entry in a sample vector $\mathbf{x}\in\mathbb{R}^{d}$ with probability $p_{mask}$, and the objective task is to predict the masked-out features from the unmasked features. Formally, let $m\in\mathbb{R}^{d}$ be a binary vector taking value $1$ when the corresponding entry in $\mathbf{x}$ is masked, $\mathbf{x}^{m}=\{x_{j}:m_{j}=1\}$ represents the masked entries of sample $\mathbf{x}$, and $\mathbf{x}^{o}=\{x_{j}:m_{j}=0\}$ its unmasked counterpart. We evaluate the benefit of replacing the traditional reconstruction learning objective defined as
(3) $\min_{\theta\in\Theta}\sum_{\mathbf{x}\in\mathcal{D}_{train}}d(\mathbf{x}^{m},\phi_{\theta}(\mathbf{x}^{o})),$
where $d(.,.)$ is a distance measure; with its equivalent augmented with a retrieval module as follows
(4) $\min_{\theta\in\Theta}\sum_{\mathbf{x}\in\mathcal{D}_{train}}d\left(\mathbf{x}^{m},\phi_{\theta}\left(\mathbf{x}^{o}\mid\mathcal{D}_{train}^{o}\right)\right),$
where $\mathcal{D}_{train}^{o}=\{\mathbf{x}_{i}^{o}\in\mathbb{R}^{d}\}_{i=1}^{n}$. In inference, $\mathcal{D}_{train}^{o}$ is replaced by $\mathcal{D}_{train}$ in which none of the features of the training samples are masked.
Table 1. Comparison of transformer-based methods.
[…]
F1-Score $56.2$ $56.1$ $55.7$ $\mathbf{58.6}$ $53.9$
AUROC $83.4$ $83.1$ $83.1$ $\mathbf{84.4}$ $82.1$
3.2. Retrieval methods
Let $\mathbf{z}$ denote the sample of interest for which we wish to reconstruct its masked features $\mathbf{z}^{m}$ given its observed counterpart $\mathbf{z}^{o}$. Let $\mathcal{C}$ denote the candidate samples from the training set from which $k$ helpers are to be retrieved, and $\mathcal{H}$ the retrieved helpers, $\mathcal{H}\subseteq\mathcal{C}$.
We consider several external retrieval modules that rely on similarity measures to identify relevant samples to augment the encoded representation of the sample of interest $\mathbf{z}$. It involves placing a retrieval module after the transformer encoder and before the output layer, as shown in figure 1. We investigate in section 5 the impact of modifying the location of the retrieval module and consider placing it before the encoder as an alternative.
For each method, the retrieval module consists in selecting the top-$k$ elements that maximize a similarity measure $\mathcal{S}(\cdot,\cdot)$ and use a value function to obtain representations of the chosen samples $\mathcal{V}(\cdot,\cdot)$ to be aggregated with sample $\mathbf{z}$.
KNN-based module
First, we consider a simple method that identifies the $k$ most relevant samples in $\mathcal{C}$ using a KNN approach. Formally, the similarity and value functions are defined as follows
(5) $\mathcal{S}(\mathbf{z},\mathbf{x})$ $=-\|\mathbf{e_{z}}-\mathbf{e_{x}}\|$
$\mathcal{V}(\mathbf{z},\mathbf{x})$ $=\mathbf{e_{x}},$
[…]
Attention-based modules
Second, we consider attention mechanisms to select $\mathcal{H}$. We consider three types of attention inspired by those proposed in (Gorishniy et al., 2023). First, the vanilla attention (later referred to as v-attention) where the score and value function used to select the retrieved samples are defined as
(6) $\mathcal{S}(\mathbf{z},\mathbf{x})$ $=W_{Q}(\mathbf{e_{z}})^{\top}W_{K}(\mathbf{e}_{x})$
$\mathcal{V}(\mathbf{z},\mathbf{x})$ $=W_{V}(\mathbf{e_{x}}).$
where $W_{Q},W_{K}$ and $W_{V}$ are learned parameters.
Second, we also consider another type of attention module, later referred to as attention-bsim, which involves replacing the score function defined in eq. 6 as follows
(7) $\mathcal{S}(\mathbf{z},\mathbf{x})$ $=-\left\|W_{K}(\mathbf{e_{z}})-W_{K}(\mathbf{e_{x}})\right\|^{2}$
$\mathcal{V}(\mathbf{z},\mathbf{x})$ $=W_{V}(\mathbf{e_{x}}).$
Third, we consider attention-bsim-bval a modification of the value function in eq. 7 as
(8) $\mathcal{S}(\mathbf{z},\mathbf{x})$ $=-\left\|W_{K}(\mathbf{e_{z}})-W_{K}(\mathbf{e_{x}})\right\|^{2}$
$\mathcal{V}(\mathbf{z},\mathbf{x})$ $=T(W_{K}(\mathbf{e_{z}})-W_{K}(\mathbf{e_{x}})),$
[…]
Aggregation
The retrieval modules necessitate aggregating the obtained retrieved representations $\mathcal{V}(\mathbf{z},\mathbf{x})$ with the representation of the sample of interest $\mathbf{z}$. We aggregate the value of the selected top-$k$ helpers, to be fed to the final layer
(9) $\tilde{\mathbf{e}}_{\mathbf{z}}=(1-\lambda)\cdot\mathbf{e_{z}}+\lambda\cdot\frac{1}{k}\sum_{x\in\mathcal{H}}\mathcal{V}(\mathbf{z},\mathbf{x}).$
where $\lambda\in[0,1)$ is a hyperparameter.
3.3. Anomaly score
We construct an anomaly score to assess whether a test sample belongs to the normal distribution or should be considered an anomaly. As a reconstruction-based method, our anomaly score is directly obtained from the optimized loss during training: the better the trained model reconstructs a sample, the more likely the sample is to be normal. Indeed, since the model has exclusively seen normal samples during training, it should be less able to reconstruct anomalies correctly since they stem from a different distribution. On the contrary, unseen normal samples should be well reconstructed. We rely on the squared $\ell_{2}$-norm of the difference between the reconstructed sample and the original sample for numerical features, while we use the cross-entropy loss function for categorical features.
We rely on a mask bank composed of $m$ $d$-dimensional masks to construct the anomaly score as proposed in (Thimonier et al., 2023a). We apply each mask to each validation sample and reconstruct the masked features to compute the reconstruction error for each mask. Thus, each validation sample is masked and reconstructed $m$ times. The anomaly score is constructed as the average reconstruction error over the $m$ masks. To construct the mask bank, we fix the maximum of features to be masked simultaneously $r$ and and construct $m=\sum_{k=1}^{r}\binom{d}{k}$ masks. Choosing deterministic masks instead of random masks to create the mask bank used for inference is beneficial for two reasons. First, since the model will reconstruct all features at least once, it increases the likelihood of identifying different types of anomalies. Indeed, anomalies that deviate from the normal distribution due to a single feature would only be identified if the corresponding mask hiding this feature would be included. Second, this approach ensures that all samples are masked identically to build the anomaly score. We investigate the impact of constructing a random mask bank instead of a deterministic mask bank in section 5.4.
We use the whole unmasked training set [^2] For large datasets, we resort to a random subsample of the training set for computational reasons. $\mathcal{C}=\mathcal{D}^{train}$ to predict the masked features of each sample for each of the $m$ masked vectors and construct the anomaly score for a validation sample $\mathbf{z}$ as
[…]
where $\mathcal{L}(\mathbf{z}^{(k)};\mathcal{D}^{train})$ designates the loss for the sample $\mathbf{z}$ with mask $k$.
4. Experiments
Datasets
We experiment on an extensive benchmark of tabular datasets following previous work (Shenkar and Wolf, 2022; Thimonier et al., 2023a). The benchmark is comprised of two datasets widely used in the anomaly detection literature, namely Arrhythmia and Thyroid, a second group of datasets, the ”Multi-dimensional point datasets”, obtained from the Outlier Detection DataSets (ODDS) [^3] http://odds.cs.stonybrook.edu/ containing 28 datasets. We also include three real-world datasets from (Han et al., 2022) to replace KDD and KDDRev that were excluded from our benchmark since they have been shown to be flawed (Silva et al., 2020): fraud, campaign, and backdoor. We display each dataset’s characteristics in table 6 in appendix A.1.
Table 2. Comparison of transformer+attention-bsim and transformer+KNN across values of $k$. Here $-1$ stands for $\mathcal{H}=\mathcal{C}$. Some values are N/A either because it is not relevant to compute (e.g. $-1$ for KNN) or when there are not enough samples in the training set for the selected value of $k$.
$k$ 00 $5$ $25$ $50$ $200$ $500$ $-1$
transformer+attention-bsim
Abalone $42.5\pm$$7.8$ $53.0\pm$$6.4$ $\mathbf{54.9}\pm$$5.4$ $\mathbf{55.0}\pm$$5.4$ $52.0\pm$$5.6$ $54.0\pm$$6.5$ $53.0\pm$$5.7$
Satellite $65.6\pm$$3.3$ $71.5\pm$$2.4$ $71.3\pm$$1.3$ $71.2\pm$$1.6$ $70.8\pm$$1.8$ $71.2\pm$$1.7$ $\mathbf{71.9}\pm$$1.5$
Lympho $88.3\pm$$7.6$ $91.7\pm$$8.3$ $\mathbf{93.3}\pm$$8.2$ $91.7\pm$$8.3$ N/A N/A $90.0\pm$$8.1$
Satimage $89.0\pm$$4.1$ $88.8\pm$$3.8$ $88.4\pm$$3.8$ $89.4\pm$$4.2$ $88.8\pm$$4.3$ $89.1\pm$$4.3$ $\mathbf{93.2}\pm$$1.7$
Thyroid $55.5\pm$$4.8$ $\mathbf{56.9}\pm$$5.3$ $55.9\pm$$5.2$ $56.3\pm$$5.2$ $56.4\pm$$5.2$ $55.9\pm$$5.6$ $55.8\pm$$6.3$
Cardio $81.0\pm$$4.1$ $81.2\pm$$1.6$ $\mathbf{81.9}\pm$$1.4$ $\mathbf{81.9}\pm$$1.4$ $\mathbf{81.9}\pm$$1.4$ $\mathbf{81.9}\pm$$1.4$ $80.6\pm$$2.4$
Ionosphere $88.1\pm$$2.8$ $89.4\pm$$5.0$ $90.2\pm$$4.5$ $89.8\pm$$4.3$ N/A N/A $\mathbf{91.7}\pm$$2.1$
mean $70.3$ $76.1$ $\mathbf{76.6}$ $76.5$ $70.0$ $70.4$ $\mathbf{76.6}$
mean std $4.9$ $4.7$ $4.3$ $4.3$ $3.7$ $3.9$ $4.0$
transformer+KNN
Abalone $42.5\pm$$7.8$ $42.5\pm$$9.5$ $42.3\pm$$7.3$ $42.3\pm$$7.9$ $43.0\pm$$7.9$ $\mathbf{43.3}\pm$$10.3$ N/A
Satellite $65.6\pm$$3.3$ $58.6\pm$$2.9$ $\mathbf{64.7}\pm$$3.8$ $63.3\pm$$2.1$ $63.3\pm$$2.6$ $51.6\pm$$3.3$ N/A
Lympho $88.3\pm$$7.6$ $89.2\pm$$7.9$ $\mathbf{90.0}\pm$$8.2$ $\mathbf{90.0}\pm$$8.2$ N/A N/A N/A
Satimage $89.0\pm$$4.1$ $86.8\pm$$0.4$ $91.7\pm$$3.5$ $\mathbf{92.2}\pm$$2.1$ $91.5\pm$$2.9$ $90.6\pm$$4.0$ N/A
Thyroid $55.5\pm$$4.8$ $55.5\pm$$4.9$ $55.2\pm$$4.2$ $\mathbf{56.1}\pm$$4.7$ $55.8\pm$$3.7$ $54.1\pm$$3.9$ N/A
Cardio $81.0\pm$$4.1$ $65.6\pm$$3.6$ $\mathbf{81.4}\pm$$1.4$ $\mathbf{81.4}\pm$$1.3$ $\mathbf{81.4}\pm$$1.3$ $\mathbf{81.3}\pm$$1.3$ N/A
Ionosphere $88.1\pm$$2.8$ $\mathbf{85.7}\pm$$3.4$ $81.8\pm$$3.0$ $81.6\pm$$2.3$ N/A N/A N/A
mean $70.3$ $69.1$ $\mathbf{72.4}$ $\mathbf{72.4}$ $67.0$ $64.2$ N/A
mean std $4.9$ $4.7$ $4.5$ $4.1$ $3.7$ $4.6$
Experimental Settings
Following previous work in the AD literature, (Zong et al., 2018; Bergman and Hoshen, 2020), we construct the training set with a random subsample of the normal samples representing 50% of the normal samples, we concatenate the 50% remaining with the entire set of anomalies to constitute the validation set. Similarly, we fix the decision threshold for the AD score such that the number of predicted anomalies equals the number of existing anomalies.
To evaluate to which extent sample-sample dependencies are relevant for anomaly detection, we compare models that attend to relations between samples to the vanilla transformer model. We compare the different methods discussed the F1-score ($\uparrow$) and AUROC ($\uparrow$) metrics following the literature. For each dataset, we report an average over $20$ runs for $20$ different seeds; we display the detailed results for all five transformer-based methods in tables 8 and 9 in appendix B.1 and report in table 1 the average F1 and AUROC.
We considered three regimes for the transformer dimensions depending on the dataset size. The transformer encoder comprises $2$ or $4$ layers with $4$ attention heads and hidden dimension $e\in\{8,16,32\}$, for smaller to larger datasets. We train the transformer with a mask probability $p_{mask}$ set to $0.25$ or $0.15$ and rely on the LAMB optimizer (You et al., 2020) with $\beta=(0.9,0.999)$ and also included a Lookahead (Zhang et al., 2019) wrapper with slow update rate $\alpha=0.5$ and $k=6$ steps between updates. We also include a dropout regularization with $p=0.1$ for attention weights and hidden layers. We ensure that during training, all samples from a batch are not masked simultaneously so that the retrieval module receives encoded representations of unmasked samples as it will in the inference stage. We considered the same transformer architecture and hyperparameters for the same dataset for each considered approach. For external retrieval modules, we chose for simplicity $\lambda=0.5$ for aggregation as detailed in eq. 9. We study in section 5.2 the effect of varying the value of $\lambda$ on the model’s performance. For the KNN module, we set $k=5$ as the cardinality of $\mathcal{H}$ since KNN-based anomaly detection methods (Ramaswamy et al., 2000) with $k=5$ have shown strong anomaly detection performance (Thimonier et al., 2023a). In contrast, for the attention modules, we set $\mathcal{H}=\mathcal{C}$ and use the attention weights to compute a weighted mean to be aggregated as in equation 9. We further discuss the choice of $k$ in section 5.1. Finally, depending on the dataset, we trained the model until the loss stopped improving for $50$ or $100$ consecutive epochs. We detail in appendix A.2.2 the varying hyperparameters between datasets. Each experiment in the present work can be replicated using the code made available on github [^4] https://github.com/hugothimonier/Making-Parametric-Anomaly-Detection-on-Tabular-Data-Non-Parametric-Again.
Results
As reported in table 1, we observe that not all retrieval modules induce a significant boost in anomaly detection performance. We observe that only the transformer augmented by the attention-bsim module perform significantly better than the vanilla transformer. In deed, augmenting the vanilla transformer with the retrieval module detailed in eq. 7 allows to increase the average F1-score by $4.3$% and AUROC by $1.2$%. This observation contradicts the results obtained for the supervised classification and regression tasks investigated in previous work (Gorishniy et al., 2023) where the module that obtains the best performance is attention-bsim-bval . However, let us mention that the attention-bsim-bval module involved in our work is not identical to the one put forward in (Gorishniy et al., 2023) since it does not involve any label.
5. Discussion
We randomly selected a subset of the dataset from the benchmark for the ablation study conducted hereafter. We use this subset of datasets in each experiment for computational reasons and to avoid cherry-picking the hyperparameters.
5.1. Choice of $k$
We investigate the impact of varying $k$, the number of helpers, for both the transformer augmented by attention-bsim and the KNN module since they are the two best-performing retrieval-augmented models. We keep hyperparameters constant and only make $k$ vary between runs. We report in table 2 the obtained results for both architectures for values of $k\in\{0,5,25,50,200,500,-1\}$, where $-1$ implies that $\mathcal{H}=\mathcal{C}$.
A noticeable trend is that both architectures obtain the best results with moderate numbers of helper, i.e. $k\in\{25,50\}$, and have worse performance for smaller values of $k$. This observation suggests that performance displayed in tables 1,8, and 9 could be improved for optimized values of $k$.
5.2. Choice of $\lambda$
We analyze the performance variation for the best-performing external retrieval module, namely transformer+attention-bsim, for varying values of $\lambda$. We compute the average F1-score over $10$ runs for each value of $\lambda$ in $\{0.1,0.2,\dots,0.7\}$ while keeping the same hyperparameters. We report the results in table 4.
We observe a slight variation of the average metric across the datasets when setting lambda values from $0.1$ to $0.7$. The maximum value is obtained for $0.5$, but the difference with other values of $\lambda$ is non-significant. Nevertheless, we observe significant differences between the obtained results for isolated datasets for different values of $\lambda$. This observation supports the idea that an optimal value of $\lambda$ exists for each dataset, which may differ between datasets.
5.3. Location of the retrieval module
Table 3. Comparison of the performance of the transformer + attention-bsim model for different retrieval module architecture based on the F1-score ($\uparrow$).
Retrieval post-enc post-emb post-emb
Aggregation post-enc post-enc post-emb
Abalone $\mathbf{53.0}\pm$$5.7$ $47.0\pm$$7.6$ $30.5\pm$$14$
Satellite $\mathbf{71.9}\pm$$1.5$ $60.5\pm$$0.9$ $65.1\pm$$2.3$
Lympho $\mathbf{90.0}\pm$$8.1$ $\mathbf{91.6}\pm$$8.3$ $84.2\pm$$11.1$
Satimage $\mathbf{93.2}\pm$$1.7$ $50.8\pm$$17.0$ $65.9\pm$$15.3$
Thyroid $55.8\pm$$6.3$ $55.2\pm$$5.9$ $\mathbf{58.1}\pm$$5.0$
Ionosphere $\mathbf{91.7}\pm$$2.1$ $88.2\pm$$1.4$ $\mathbf{91.3}\pm$$2.0$
mean $\mathbf{75.9}$ $65.6$ $65.9$
mean std $\mathbf{4.2}$ $6.9$ $8.3$
We investigate the impact of the retrieval module’s location on the retrieval-augmented models’ performance. To do so, we focus on the transformer+attention-bsim model since it has shown to be the best-performing retrieval-based method. We compare three different architectures:
  • (post-enc, post-enc) for post-encoder location and post-encoder aggregation: the architecture detailed in figure 1,
  • (post-emb, post-enc), the architecture in which the retrieval module is located after the embedding layer, but the aggregation is still located after the encoder,
  • (post-emb, post-emb) the architecture where both retrieval and aggregation are located after the embedding layer.
[…]
We observe that the (post-enc, post-enc) architecture obtains the highest mean and lowest mean standard deviation over the tested datasets by a sizable margin. The transformer encoder’s expressiveness allows for better representations of a data sample than the embedding layer and may account for such results. Indeed, this shows that the retrieval modules that receive the embedded representation as inputs are less able to select the relevant sample to foster mask reconstruction and anomaly detection performance. Moreover, since the encoder and retrieval modules are trained conjointly, the retrieval module can help the encoder converge to a state that favors sample representations that allow relevant clusters to be formed.
Table 4. Comparison of transformer+attention-bsim across values of $\lambda$
$\lambda$ 00 $0.1$ $0.2$ $0.3$ $0.4$ $0.5$ $0.6$ $0.7$
Abalone $42.5$ $44.3$ $47$ $47.3$ $46.5$ $\mathbf{53}$ $46.5$ $45.8$
Satellite $65.6$ $72$ $72$ $72.5$ $73$ $71.9$ $\mathbf{73.5}$ $73.4$
Lympho $88.3$ $90.8$ $90.8$ $91.7$ $\mathbf{92.5}$ $90$ $91.7$ $90$
Satim. $89$ $94.5$ $\mathbf{94.6}$ $94.1$ $94.1$ $93.2$ $93.8$ $94.1$
Thyroid $55.5$ $57.2$ $\mathbf{57.4}$ $56.6$ $56.6$ $55.8$ $57.1$ $57.3$
Cardio $68.8$ $\mathbf{80.9}$ $80.6$ $\mathbf{80.9}$ $\mathbf{80.9}$ $\mathbf{80.9}$ $\mathbf{80.9}$ $80.7$
Ionosp. $88.1$ $90.3$ $92.1$ $92.2$ $91.2$ $91.7$ $\mathbf{92.3}$ $91.7$
mean $71.1$ $75.7$ $76.4$ $76.5$ $76.4$ $\mathbf{76.6}$ $76.5$ $76.1$
5.4. Random mask bank
We also investigate the impact of constructing a random mask bank for inference instead of selecting a deterministic mask bank as discussed in section 3.3. We construct for inference a random mask bank composed of the same number of masks as for the deterministic mask bank and use the same probability $p_{mask}$ as used to train the model. We compare the performance of the transformer model+attention-bsim based on the F1-score for the two set-ups and display the results in table 5. We observe that the deterministic mask bank detects anomalies better on most tested datasets. When computing the anomaly score with the deterministic mask bank, the model obtains an average F1-score of $76.6$ over the $7$ datasets, while with the random mask bank, the model obtains $65.4$. Moreover, as expected, we also observed a significantly larger standard deviation between runs. We might expect the standard deviation to decrease as the number of masks increases, which would induce significant computational overhead.
Table 5. Comparison of the performance of the transformer + attention-bsim model for two mask bank set-ups. The same model was used for inference in both set-ups. We report an average over $10$ different splits of the data.
Mask bank random deterministic
Abalone $43.0\pm$$14.9$ $\mathbf{53.0}\pm$$5.7$
Satellite $58.4\pm$$5.5$ $\mathbf{71.9}\pm$$1.5$
Lympho $86.7\pm$$8.5$ $\mathbf{90.0}\pm$$8.1$
Satimage $47.7\pm$$20.5$ $\mathbf{93.2}\pm$$1.7$
Thyroid $\mathbf{55.6}\pm$$6.1$ $\mathbf{55.8}\pm$$6.3$
Cardio $\mathbf{81.3}\pm$$1.6$ $80.6\pm$$2.4$
Ionosphere $85.2\pm$$4.3$ $\mathbf{91.7}\pm$$2.1$
mean $65.4$ $\mathbf{76.6}$
mean std $8.8$ $\mathbf{4.0}$
6. Limitations and Conclusion
Limitations
As with most non-parametric models, our retrieval-augmented models display a higher complexity than parametric approaches. These approaches can scale well for datasets with a reasonable number of features $d$; however, for large values of $d$, these models involve a high memory cost.
Conclusion
In this work, we have proposed an extensive investigation into external retrieval to augment reconstruction-based anomaly detection methods for tabular data. We have shown that augmenting existing AD methods using attention-based retrieval modules can help foster performance by allowing the model to attend to sample-sample dependencies. Indeed, our experiments involving an extensive benchmark of tabular datasets demonstrate the effectiveness of retrieval-based approaches since the architecture involving the attention-bsim module surpasses the vanilla transformer by a significant margin. Overall, this work emphasizes the pertinence of leveraging sample-sample dependencies to detect anomalies on tabular datasets.
Future Work
Overall, our work showed that the best-performing attention-based retrieval mechanism relies on other forms of attention than vanilla attention. Parallel to that, models like NPT-AD Thimonier et al. (2023a) have shown strong performance for anomaly detection on tabular data and rely on standard multi-head self-attention through the Attention Between Datapoints (ABD) mechanism, close to the v-attention module, to leverage inter-sample relations. Our findings may invite further research on modifying the ABD mechanism involved in NPTs to improve their AD performance. Finally, the use of external retrieval modules proved effective for the task of anomaly detection using mask reconstruction. The proposed external retrieval module could also be easily added to existing deep-AD methods to test whether they may prove relevant for other pretext tasks for anomaly detection on tabular data.
Appendix A Datasets characteristics and experimental settings
A.1. Dataset characteristics
In table 6, we display the main characteristics of the datasets involved in our experiments.
Table 6. Datasets characteristics
Dataset $n$ $d$ Outliers
Wine $129$ $13$ $10$ ($7.7$%)
Lympho $148$ $18$ $6$ ($4.1$%)
Glass $214$ $9$ $9$ ($4.2$%)
Vertebral $240$ $6$ $30$ ($12.5$%)
WBC $278$ $30$ $21$ ($5.6$%)
Ecoli $336$ $7$ $9$ ($2.6$%)
Ionosphere $351$ $33$ $126$ ($36$%)
Arrhythmia $452$ $274$ $66$ ($15$%)
BreastW $683$ $9$ $239$ ($35$%)
Pima $768$ $8$ $268$ ($35$%)
Vowels $1456$ $12$ $50$ ($3.4$%)
Letter Recognition $1600$ $32$ $100$ ($6.25$%)
Cardio $1831$ $21$ $176$ ($9.6$%)
Seismic $2584$ $11$ $170$ ($6.5$%)
Musk $3062$ $166$ $97$ ($3.2$%)
Speech $3686$ $400$ $61$ ($1.65$%)
Thyroid $3772$ $6$ $93$ ($2.5$%)
Abalone $4177$ $9$ $29$ ($0.69$%)
Optdigits $5216$ $64$ $150$ ($3$%)
Satimage-2 $5803$ $36$ $71$ ($1.2$%)
Satellite $6435$ $36$ $2036$ ($32$%)
Pendigits $6870$ $16$ $156$ ($2.27$%)
Annthyroid $7200$ $6$ $534$ ($7.42$%)
Mnist $7603$ $100$ $700$ ($9.2$%)
Mammography $11183$ $6$ $260$ ($2.32$%)
Shuttle $49097$ $9$ $3511$ ($7$%)
Mulcross $262144$ $4$ $26214$ ($10$%)
ForestCover $286048$ $10$ $2747$ ($0.9$%)
Campaign $41188$ $62$ $4640$ ($11.3$%)
Fraud $284807$ $29$ $492$ ($0.17$%)
Backdoor $95329$ $196$ $2329$ ($2.44$%)
A.2. Experimental settings
A.2.1. Training pipeline
Let $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ be a sample with $d$ features, which can be either numerical or categorical. Let $e$ designate the hidden dimension of the transformer. The training pipeline consists of the following steps:
  • Masking: We sample from a Bernouilli distribution with probability $p_{mask}$ whether each of the $d$ features is masked.
    $\mbox{mask}=(m_{1},\dots,m_{d}),$
    where $m_{j}\sim\mathcal{B}(1,p_{mask})$ $\forall j\in[1,...,d]$ and $m_{j}=1$ if feature $j$ is masked.
  • Encoding: For numerical features, we normalize to obtain 00 mean and unit variance, while we use one-hot encoding for categorical features. At this point, each feature $j$ for $j\in[1,2,...,d]$ has an $e_{j}$-dimensional representation, $encoded(\mathbf{x}_{j})\in\mathbb{R}^{e_{j}}$ , where $e_{j}=1$ for numerical features and for categorical features $e_{j}$ corresponds to its cardinality. We then mask each feature according to the sampled mask vector and concatenate each feature representation with the corresponding mask indicator function. Hence, each feature $j$ has an $(e_{j}+1)$-dimensional representation
    $((1-m_{j})\cdot encoded(\mathbf{x}_{j}),m_{j})\in\mathbb{R}^{e_{j}+1},$
    where $\mathbf{x}_{j}$ is the $j$-th features of sample $\mathbf{x}$.
  • In-Embedding: We pass each of the features encoded representations of sample $\mathbf{x}$ through learned linear layers $\texttt{Linear}(e_{j}+1,e)$. We also learn $e$-dimensional index and feature-type embeddings as proposed in (Kossen et al., 2021). Both are added to the embedded representation of sample $\mathbf{x}$. The obtained embedded representation is thus of dimension $d\times e$
    $e_{\mathbf{x}}=(e_{\mathbf{x}_{1}},e_{\mathbf{x}_{2}},\dots,e_{\mathbf{x}_{d}})\in\mathbb{R}^{d\times e},$
    and $e_{\mathbf{x}_{1}}\in\mathbb{R}^{e}$.
  • Transformer Encoder: The embedded representation obtained from the previous step is then passed through a transformer encoder. The output of the transformer $e_{\mathbf{x}}$ is of dimension $d\times e$.
  • Out-Embedding: The output of the transformer, $e_{\mathbf{x}}\in\mathbb{R}^{d\times e}$ is then used to compute an estimation of the original $d$-dimensional vector. To obtain the estimated feature $j$, we take the $j$-th $d$-dimensional representation which is output by the transformer encoder, $e_{\mathbf{x}_{j}}\in\mathbb{R}^{d}$, and pass it through a linear layer $\texttt{Linear}(e,e_{j})$, where $e_{j}$ is $1$ for numerical features and the cardinality for categorical features.
  • External Retrieval Modules: During training, for a batch $\mathcal{B}$ composed of $b$ samples, for each sample $\mathbf{x}\in\mathcal{B}$, the entire batch serves as the candidates $\mathcal{C}$. In inference, a random subsample of the training set is used as $\mathcal{C}$. Both for training and inference, when possible memory-wise, we use as $\mathcal{B}$ and $\mathcal{C}$ the entire training set.
    As input, the retrieval module receives a $\mathbb{R}^{d\times e}$ representation for each sample. Operations described in eq. 5, 6, 7 and 8 are performed on the flattened representations of samples $\mathbf{x}_{flatten}\in\mathbb{R}^{d\cdot e}$. After selecting $\mathcal{H}\subseteq\mathcal{C}$, each sample is transformed back to its original dimension to allow aggregation as described in equation 9.
A.2.2. Hyperparameters
For the vanilla transformer, as discussed in section 4, we chose a transformer model comprised of either $2$ or $4$ layers with $4$ attention heads for each model, whether it includes a retrieval module or not. We display in table 7 the varying parameters between datasets for the transformer model. Let us mention that minor changes between retrieval modules may have been included to account for memory requirements (especially for the attention-bsim-bval retrieval module, which we found to be more memory costly). We implemented early stopping with various number of epochs depending on the datasets, we refer the reader to the code made available online for more detail on each run. This lead to a different number of epoch for different seed, we thus do not display in the following table the number of epochs. Overall, learning rate, masking probability, $r$ the maximum number of features masked simultaneously, and $e$ the hidden dimension was kept unchanged across architecture.
Table 7. Datasets hyperparameters for the transformer-based approaches. When the batch size is $-1$ it refers to a full pass over the training set before an update of the weights. Due to memory requirements, for some retrieval-based approaches, the batch size may have been adjusted. The column card$(\mathcal{C})$ corresponds to the cardinality of the random subsamples of the training set used to select the $\mathcal{H}$ during inference. During training, $\mathcal{C}$ corresponds to the batch and has the batch size as cardinality. We refer the reader to the code made available online for more details on the hyperparameters.
Dataset num layers batch size lr $p^{train}_{mask}$ $r$ $m$ $e$ card$(\mathcal{C})$
Wine $2$ $-1$ $0.001$ $0.15$ $3$ $78$ $8$ $-1$
Lympho $2$ $-1$ $0.001$ $0.15$ $1$ $18$ $16$ $-1$
Glass $2$ $-1$ $0.001$ $0.15$ $1$ $9$ $16$ $-1$
Vertebral $2$ $-1$ $0.001$ $0.15$ $1$ $6$ $8$ $-1$
WBC $2$ $-1$ $0.01$ $0.15$ $1$ $30$ $16$ $-1$
Ecoli $2$ $-1$ $0.001$ $0.15$ $1$ $7$ $16$ $-1$
Ionosphere $2$ $-1$ $0.001$ $0.15$ $1$ $33$ $8$ $-1$
Arrhythmia $2$ $-1$ $0.001$ $0.15$ $1$ $274$ $8$ $-1$
BreastW $4$ $-1$ $0.001$ $0.15$ $1$ $9$ $8$ $-1$
Pima $4$ $-1$ $0.001$ $0.15$ $4$ $162$ $8$ $-1$
Vowels $4$ $-1$ $0.001$ $0.15$ $2$ $78$ $16$ $-1$
Letter Recognition $4$ $-1$ $0.001$ $0.25$ $1$ $32$ $16$ $-1$
Cardio $4$ $-1$ $0.001$ $0.15$ $2$ $231$ $16$ $-1$
Seismic $4$ $-1$ $0.001$ $0.25$ $1$ $11$ $16$ $-1$
Musk $4$ $-1$ $0.001$ $0.15$ $1$ $166$ $8$ $-1$
Speech $4$ $128$ $0.001$ $0.15$ $1$ $400$ $8$ $64$
Thyroid $4$ $-1$ $0.001$ $0.15$ $3$ $41$ $8$ $-1$
Abalone $4$ $-1$ $0.001$ $0.15$ $2$ $45$ $16$ $-1$
Optdigits $4$ $2048$ $0.001$ $0.15$ $1$ $64$ $16$ $1024$
Satimage-2 $4$ $2048$ $0.001$ $0.2$ $1$ $36$ $16$ $1024$
Satellite $4$ $2048$ $0.001$ $0.15$ $1$ $36$ $16$ $1024$
Pendigits $4$ $-1$ $0.001$ $0.15$ $1$ $16$ $16$ $-1$
Annthyroid $4$ $-1$ $0.001$ $0.15$ $1$ $6$ $16$ $-1$
Mnist $4$ $-1$ $0.001$ $0.15$ $1$ $100$ $8$ $256$
Mammography $4$ $-1$ $0.001$ $0.15$ $1$ $6$ $16$ $1024$
Shuttle $4$ $2048$ $0.001$ $0.15$ $3$ $129$ $16$ $1024$
Mulcross $4$ $2048$ $0.001$ $0.15$ $1$ $4$ $16$ $1024$
ForestCover $4$ $1024$ $0.001$ $0.15$ $2$ $55$ $16$ $1024$
Campaign $4$ $2048$ $0.001$ $0.15$ $1$ $62$ $16$ $512$
Fraud $4$ $4096$ $0.001$ $0.2$ $1$ $29$ $32$ $1024$
Backdoor $4$ $256$ $0.001$ $0.2$ $1$ $196$ $32$ $512$
Appendix B Detailed experiments
B.1. Detailed tables for main experiments
Table 8. Anomaly detection F1-score ($\uparrow$). We perform $5$% T-test to test whether the difference between the highest metrics for each dataset is statistically significant.
[…]
-bval
Wine $23.5\pm$$7.9$ $24.9\pm$$5.9$ $26.5\pm$$7.3$ $\mathbf{29.0}\pm$$7.7$ $27.0\pm$$5.6$
Lympho $88.3\pm$$7.6$ $89.2\pm$$7.9$ $88.3\pm$$9.3$ $\mathbf{90.0}\pm$$8.1$ $89.1\pm$$7.9$
Glass $\mathbf{14.4}\pm$$6.1$ $12.8\pm$$3.9$ $12.2\pm$$3.3$ $12.8\pm$$3.9$ $11.1\pm$$0.1$
Verteb. $12.3\pm$$5.2$ $\mathbf{15.7}\pm$$2.8$ $12.7\pm$$4.1$ $14.3\pm$$2.6$ $13.5\pm$$5.1$
Wbc $66.4\pm$$3.2$ $65.2\pm$$4.0$ $66.2\pm$$5.2$ $65.5\pm$$4.4$ $\mathbf{67.9}\pm$$4.2$
Ecoli $75.0\pm$$9.9$ $\mathbf{75.6}\pm$$7.5$ $\mathbf{75.6}\pm$$9.7$ $73.8\pm$$8.0$ $75.0\pm$$9.7$
Ionosph. $88.1\pm$$2.8$ $85.7\pm$$3.4$ $86.0\pm$$4.7$ $\mathbf{91.7}\pm$$2.1$ $79.3\pm$$1.6$
Arrhyth. $59.8\pm$$2.2$ $60.3\pm$$2.2$ $60.2\pm$$2.7$ $\mathbf{61.2}\pm$$2.1$ $\mathbf{61.1}\pm$$2.8$
Breastw $96.7\pm$$0.3$ $96.7\pm$$0.3$ $\mathbf{96.8}\pm$$0.3$ $96.7\pm$$0.3$ $96.7\pm$$0.3$
Pima $65.6\pm$$2.0$ $64.7\pm$$3.1$ $64.0\pm$$3.3$ $64.3\pm$$2.4$ $\mathbf{67.0}\pm$$1.5$
Vowels $28.7\pm$$8.0$ $40.0\pm$$10.0$ $49.1\pm$$11.1$ $44.5\pm$$10.5$ $\mathbf{58.0}\pm$$11.2$
Letter $41.5\pm$$6.2$ $32.9\pm$$11.8$ $41.8\pm$$11.5$ $\mathbf{43.7}\pm$$10.3$ $28.5\pm$$7.1$
Cardio $\mathbf{68.8}\pm$$2.8$ $65.6\pm$$3.6$ $62.8\pm$$6.9$ $67.7\pm$$3.7$ $68.3\pm$$4.5$
Seismic $\mathbf{19.1}\pm$$5.7$ $17.4\pm$$5.5$ $\mathbf{19.5}\pm$$6.3$ $16.7\pm$$5.5$ $17.5\pm$$5.4$
Musk $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$
Speech $\mathbf{6.8}\pm$$1.9$ $6.3\pm$$1.4$ $5.7\pm$$1.7$ $5.9\pm$$1.5$ $5.9\pm$$1.7$
Thyroid $55.5\pm$$4.8$ $55.5\pm$$4.9$ $\mathbf{56.0}\pm$$5.9$ $\mathbf{55.8}\pm$$6.3$ $55.3\pm$$6.6$
Abalone $42.5\pm$$7.8$ $42.5\pm$$9.5$ $49.8\pm$$6.5$ $\mathbf{53.0}\pm$$5.7$ $43.2\pm$$9.1$
Optdig. $61.1\pm$$4.7$ $\mathbf{70.7}\pm$$16.5$ $51.5\pm$$7.6$ $62.6\pm$$6.5$ $22.8\pm$$5$
Satimage2 $89.0\pm$$4.1$ $86.8\pm$$0.4$ $90.7\pm$$2.6$ $\mathbf{93.2}\pm$$1.7$ $64.2\pm$$7.2$
Satellite $65.6\pm$$3.3$ $58.6\pm$$2.9$ $57.3\pm$$3.0$ $\mathbf{71.9}\pm$$1.5$ $53.7\pm$$3.3$
Pendig. $35.4\pm$$10.9$ $\mathbf{52.1}\pm$$9.0$ $39.0\pm$$14.5$ $\mathbf{53.4}\pm$$9.8$ $34.2\pm$$12.2$
Annthyr. $29.9\pm$$1.5$ $\mathbf{30.4}\pm$$1.9$ $\mathbf{30.3}\pm$$1.5$ $\mathbf{30.3}\pm$$1.6$ $\mathbf{30.5}\pm$$1.4$
Mnist $56.7\pm$$5.7$ $\mathbf{64.2}\pm$$3.7$ $61.7\pm$$1.0$ $61.6\pm$$1.0$ $56.7\pm$$1.9$
Mammo. $17.4\pm$$2.2$ $17.3\pm$$2.4$ $15.5\pm$$2.5$ $17.2\pm$$3.0$ $\mathbf{17.7}\pm$$2.7$
Shuttle $85.3\pm$$9.8$ $90.8\pm$$2.9$ $67.7\pm$$13.7$ $87.8\pm$$3.7$ $\mathbf{95.6}\pm$$1.8$
Mullcr. $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$
Forest $21.3\pm$$3.1$ $18.6\pm$$4.6$ $21.0\pm$$5.9$ $\mathbf{24.9}\pm$$6.5$ $11.1\pm$$4.1$
Camp. $47.0\pm$$1.9$ $48.5\pm$$2.1$ $43.3\pm$$2.3$ $49.7\pm$$1.2$ $49.1\pm$$1.1$
Fraud $53.4\pm$$4.4$ $56.4\pm$$2,1$ $56.3\pm$$2.1$ $57.1\pm$$2.1$ $55.2\pm$$1.8$
Backd. $85.8\pm$$0.6$ $\mathbf{86.1}\pm$ $0.6$ $85.2\pm$ $0.7$ $85.3\pm$ $0.6$ $82.4\pm$ $1.3$
mean $56.2$ $56.1$ $55.7$ $\mathbf{58.6}$ $53.9$
mean std $4.4$ $4.6$ $5.0$ $\mathbf{3.7}$ $\mathbf{3.7}$
Table 9. Anomaly detection AUROC($\uparrow$). We perform $5$% T-test to test whether the differences between the highest metrics for each dataset are statistically significant.
[…]
-bval
Wine $61.4\pm$$6.7$ $60.4\pm$$5.4$ $62.1\pm$$6.4$ $\mathbf{63.5}\pm$$7.8$ $\mathbf{64.5}\pm$$5.6$
Lympho $99.5\pm$$0.4$ $99.6\pm$$0.4$ $99.6\pm$$0.5$ $\mathbf{99.7}\pm$$0.3$ $\mathbf{99.7}\pm$$0.3$
Glass $61.2\pm$$7.0$ $61.2\pm$$5.0$ $\mathbf{62.1}\pm$$7.0$ $59.3\pm$$6.9$ $59.1\pm$$5.8$
Vertebral $44.8\pm$$5.2$ $\mathbf{46.7}\pm$$4.1$ $45.3\pm$$7.1$ $45.4\pm$$3.7$ $45.4\pm$$4.7$
WBC $95.0\pm$$1.1$ $94.3\pm$$1.5$ $94.3\pm$$1.6$ $94.2\pm$$1.1$ $\mathbf{95.5}\pm$$1.6$
Ecoli $84.8\pm$$1.6$ $84.8\pm$$1.8$ $85.2\pm$$2.7$ $\mathbf{87.4}\pm$$1.8$ $85.4\pm$$2.3$
Ionosph. $95.4\pm$$1.9$ $93.7\pm$$2.7$ $93.6\pm$$4.0$ $\mathbf{97.5}\pm$$0.1$ $87.2\pm$$2.3$
Arrhyth. $81.7\pm$$1.1$ $81.9\pm$$0.9$ $81.8\pm$$0.9$ $\mathbf{82.3}\pm$$0.7$ $\mathbf{82.1}\pm$$0.9$
Breastw $\mathbf{99.6}\pm$$0.1$ $\mathbf{99.6}\pm$$0.1$ $\mathbf{99.6}\pm$$0.1$ $\mathbf{99.6}\pm$$0.1$ $\mathbf{99.6}\pm$$0.1$
Pima $67.2\pm$$2.4$ $66.0\pm$$3.8$ $65.4\pm$$3.8$ $65.8\pm$$2.9$ $\mathbf{68.7}\pm$$1.4$
Vowels $78.4\pm$$9.2$ $86.1\pm$$5.2$ $90.4\pm$$4.7$ $88.3\pm$$4.5$ $\mathbf{94.3}\pm$$2.8$
Letter $80.5\pm$$4.8$ $73.5\pm$$9.6$ $\mathbf{81.0}\pm$$8.7$ $\mathbf{81.5}\pm$$6.8$ $69.1\pm$$7.7$
Cardio $\mathbf{93.5}\pm$$1.3$ $92.0\pm$$1.7$ $89.9\pm$$4.2$ $\mathbf{93.3}\pm$$1.7$ $\mathbf{93.7}\pm$$1.3$
Seismic $\mathbf{58.2}\pm$$7.9$ $56.8\pm$$8.4$ $57.9\pm$$7.6$ $\mathbf{58.0}\pm$$6.7$ $54.8\pm$$6.2$
Musk $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$
Speech $47.2\pm$$0.7$ $\mathbf{47.3}\pm$$0.8$ $\mathbf{47.3}\pm$$0.8$ $\mathbf{47.3}\pm$$0.8$ $47.0\pm$$0.5$
Thyroid $\mathbf{93.8}\pm$$1.2$ $\mathbf{93.8}\pm$$1.2$ $93.6\pm$$5.9$ $93.7\pm$$1.5$ $93.6\pm$$1.8$
Abalone $\mathbf{88.3}\pm$$2.0$ $86.1\pm$$3.6$ $\mathbf{88.0}\pm$$3.5$ $\mathbf{87.9}\pm$$3.7$ $86.6\pm$$2.9$
Optdig. $96.4\pm$$4.7$ $96.2\pm$$9.8$ $94.9\pm$$1.7$ $\mathbf{96.7}\pm$$1.1$ $83.4\pm$$3.2$
Satimage $99.7\pm$$0.1$ $99.5\pm$$0.2$ $99.6\pm$$0.2$ $\mathbf{99.8}\pm$$0.1$ $96.8\pm$$1.9$
Satellite $73.8\pm$$2.5$ $68.9\pm$$2.0$ $67.8\pm$$2.5$ $\mathbf{79.5}\pm$$1.9$ $62.0\pm$$2.9$
Pendigits $93.8\pm$$2.6$ $96.5\pm$$1.4$ $94.1\pm$$2.8$ $\mathbf{97.1}\pm$$1.2$ $89.4\pm$$7.1$
Annthyr. $65.4\pm$$1.4$ $66.0\pm$$1.7$ $\mathbf{66.2}\pm$$1.3$ $\mathbf{66.2}\pm$$1.6$ $66.0\pm$$1.1$
Mnist $87.4\pm$$3.2$ $\mathbf{90.3}\pm$$2.2$ $89.9\pm$$0.4$ $\mathbf{90.0}\pm$$0.5$ $87.3\pm$$1.0$
Mammo. $77.6\pm$$1.0$ $76.8\pm$$2.4$ $75.2\pm$$2.9$ $78.4\pm$$1.6$ $\mathbf{79.8}\pm$$1.8$
Mullcross $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$ $\mathbf{100}\pm$$0.0$
Shuttle $97.2\pm$$2.2$ $98.1\pm$$0.5$ $90.2\pm$$6.1$ $97.7\pm$$0.9$ $\mathbf{98.9}\pm$$0.3$
Forest $95.1\pm$$0.8$ $94.5\pm$$0.7$ $95.0\pm$$1.0$ $\mathbf{95.4}\pm$$0.9$ $92.7\pm$$0.7$
Campaign $75.3\pm$$2.1$ $75.7\pm$$1.9$ $69.3\pm$$2.0$ $76.1\pm$$2.0$ $75.9\pm$$2.1$
Fraud $94.7\pm$$0.4$ $95.1\pm$$0.4$ $95.2\pm$$0.4$ $\mathbf{95.8}\pm$$0.4$ $94.7\pm$$0.4$
Backdoor $\mathbf{95.1}\pm$$0.2$ $\mathbf{95.2}\pm$$0.3$ $94.5\pm$$0.2$ $94.7\pm$$0.1$ $91.7\pm$$0.3$
mean $83.4$ $83.1$ $83.1$ $\mathbf{84.4}$ $82.1$
mean std $2.4$ $2.8$ $2.6$ $2.0$ $2.3$
B.2. Comparison to existing models
(a) F1-score ($\uparrow$)
(b) Rank ($\downarrow$)
Figure 2. For each of the $31$ datasets on which models were evaluated,we report the average F1-score over $20$ runs for $20$ different seeds. See (Thimonier et al., 2023a) for detail on the obtained metrics. For both figures, the model displayed on the far left is the worst-performing model for the chosen metric, and on the far right is the best-performing model. We also highlight the metric of the best-performing model in bold.
For completeness we compare the different architectures proposed in the present work to existing methods in the literature. We rely on the experiments conducted in (Thimonier et al., 2023a) and Shenkar and Wolf (2022) for the metrics of the competing methods. We display in figure 2 the comparison to existing methods. We compare our methods to recent deep methods, namely GOAD (Bergman and Hoshen, 2020), DROCC (Goyal et al., 2020), NeuTraL-AD (Qiu et al., 2021), the contrastive approach proposed in (Shenkar and Wolf, 2022) and NPT-AD (Thimonier et al., 2023a). We also compare to classical non-deep methods such as Isolation Forest (Liu et al., 2008), KNN (Ramaswamy et al., 2000), RRCF (Guha et al., 2016), COPOD (Li et al., 2020) and PIDForest (Gopalan et al., 2019). We refer the reader to Thimonier et al. (2023a) for the detail on the F1-score per dataset for other methods than those shown in tables 8 and 9.